Circular No. 6682

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cr IN ANONYMOUS GALAXY M. Turatto and S. Benetti, European Southern Observatory; and E.

Cappellaro and A. Piemonte, Padova, report: “A low-resolution spectrogram (range 500–900 nm, resolution 1.2 nm) of SN 1997cr (IAUC 6679) was obtained under poor conditions on June 14.42 UT at the Danish 1.5-m telescope (+ DFOSC) at La Silla. The low-signal-to-noise spectrum shows a blue continuum and only one broad (FWHM ∼ 25 nm) emission centered at 707 nm, with no P-Cyg absorption. When this feature is identified with Hα, the spectrum matches that of the bright linear type-II SN 1980K around maximum and implies a redshift z = 0.077 for the parent galaxy. The absolute magnitude (M_{V} ≥ −17, H = 75 km s^{−1} Mpc^{−1}; cf. IAUC 6679) is consistent with classification of SN 1997cr as type II. It must be noted, however, that if this interpretation is correct, the parent galaxy does not belong to Abell 3744 itself, but rather to one of the background galaxies detected in the survey by Katgert et al. (1996, A.Ap. 310, 8).”

COMET C/1995 O1 (HALE-BOPP) N. Dello Russo, M. J. Mumma, and M. A. DiSanti, Goddard Space

Flight Center, NASA; K. Magee-Sauer, Rowan University; R. Novak, Iona College; and T. Rettig, University of Notre Dame, report: “We detected infrared emission from OCS and CN in C/1995 O1, using the NASA Infrared Telescope Facility (+ CSHELL cryogenic infrared spectrometer) on Mauna Kea. Lines (P3, P10, and R3) of the CN 1-0 vibrational band were detected on seven occasions during the interval Mar. 1.9–May 1.2 UT. Nine lines (P10 to P24) of the OCS ν_{3} vibrational band were detected near 4.86 μm. For a rotational temperature of 100 K, the production rate (× 10^{27} molecules/s) of OCS was 3.4 on Apr. 16.2, and 3.2 on Apr. 30.1 and May 1.2. The water-production rate was 5 × 10^{30}, based on our measurements of lines of the hot band (100–010). The mixing ratio is then Q(OCS)/Q(H_{2}O) = 7 × 10^{−4}. The OCS lines were sharply peaked on the nucleus, while the CN lines were extended over the length of the slit (± 15′′ about the nucleus), as expected.”

NOVA SCORPII 1997 Visual magnitude estimates: June 10.20 UT, 10.4 (W. G. Dillon, Mis-

souri City, TX); 12.10, 11.7 (A. Pereira, Cabo da Roca, Portugal); 14.99, 11.2 (F. Pujol, Madrid, Spain); 14.99, 11.3 (D. Rodriguez, Madrid, Spain).

1997 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green